Three isotopes of an element have mass numbers M,(M+1),(M+2).If the mean mass number is(M+0.5),then what mole ratio may be accepted for them in order.
Please give the solution.

Grade:10

1 Answers

Sunil Kumar FP
10 years ago
acoording to your question
let the relative abundundance in decimal be x y and z for M,M+1,M+2 respectively
then x+y+z=1
given
Mx +(M+1)y +(M+2)z=M+.5
y+2z=.5
no other thing is given and there can be various ratio possible base on only 1eqn.as the average relative mass is close to M therefore it should have large mole of ratio.

thanks and regards
sunil kr
askIItian faculty